OpenStudy (theedwardsfamily):

Find the mean of this set of numbers. 5, –11, 7, 12, –1, 0 A. –6 B. –2 C. 2 D. 6

OpenStudy (theedwardsfamily):

@texaschic101 !

myininaya (myininaya):

add up the numbers first

OpenStudy (theedwardsfamily):

@Michele_Laino

myininaya (myininaya):

take that sum and divide it by the number of numbers you have second

OpenStudy (theedwardsfamily):

is it 2?

myininaya (myininaya):

i will check what did you get for the sum of the numbers?

OpenStudy (anonymous):

The sum of the numbers is 12, then you would divide by 6, so ya 2

OpenStudy (theedwardsfamily):

i got 12

myininaya (myininaya):

great and there are 6 numbers

myininaya (myininaya):

so 12/6=2

OpenStudy (theedwardsfamily):

im right yay

myininaya (myininaya):

very good @TheEdwardsFamily

OpenStudy (theedwardsfamily):

now who should I give the medal to @myininaya or @mememe1 ?

OpenStudy (michele_laino):

please you have to perform this summation: \[5+(-11)+7+12+(-1)+0= ...\]

OpenStudy (theedwardsfamily):

I did that and I worked it out and got 2

OpenStudy (michele_laino):

ok! now divide your result by 6

OpenStudy (michele_laino):

wait, you should get 12 as sum

OpenStudy (michele_laino):

so 12/6=2 so C is it clear, please?

OpenStudy (theedwardsfamily):

yea did that

OpenStudy (theedwardsfamily):

can you give me a medal please?

OpenStudy (michele_laino):

ok!

OpenStudy (theedwardsfamily):

thanks